OpenStudy (anonymous):

Tom has a bag of red,green and yellow sweets. He picks a sweet out of the bag without looking. the probability of him picking a red sweet is 1/7 the probability of him picking a green sweet is twice the probability of him picking a yellow sweet work out the probability of him picking a yellow sweet

OpenStudy (nubeer):

probability of yellow =x probability of green =2x probability of red = 1/7 so 1=probability of yellow +probability of green + probability of red.

OpenStudy (anonymous):

oh i see ok thanks :)

OpenStudy (nubeer):

you're welcome :)

OpenStudy (anonymous):

ok got a second part for you

OpenStudy (anonymous):

tom know there are no more than 10 sweets in the bag? how many sweets are in the bag explain your answer?

OpenStudy (anonymous):

very simple question ....answers 1/7, 2/7 , 4/7

OpenStudy (anonymous):

great job! ;) thanks

OpenStudy (anonymous):

but how do you explain it?

OpenStudy (anonymous):

1/7 + x + 2x = 1

OpenStudy (anonymous):

oh i see thanks

OpenStudy (anonymous):

does anyone else have a different method?